Show that if S and T are linear

transformations on a finite dimensional

vector space, then rank (ST)<= rank (S).


1
Expert's answer
2021-07-28T13:46:33-0400

Let , "T:V\rightarrow V" and "S:V\rightarrow V" be two linear transformation from a finite dimensional vector space "V" to itself.

Therefore, "S\circ T :V \rightarrow V" is a linear transformation. Also note,

"dim(ImT)=rank(T),

\\dim(ImS)=rank(S)"

and "dim(Im(ST))=dim(Im(S\circ T))=rank (ST)."

"dim(Ker (S))=nullity (S)\\
dim(ker(T))=nullity(T)"

and "dim(ker(ST))=nullity (ST)" .

Note: Generally,"\ S\circ T" written as "ST" .

Now ,"Ker(ST)=\{x\in V :(ST)(x)=0 \}" .

Where  is a zero vector of "V."

"=\{ x\in V:S(T(x))=0 \}"

Now , if "x\in ker(T)" then

"T(x)=0 \ \implies S(T(x))=S(0)=0".

Hence, "x\in Ker(ST)."

Therefore,"Ker(T)\sube Ker(ST)."

Thus,"dim(Ker(T))\leq dim(Ker(ST))"

I,e,"nullity (T)\leq nullity (ST)...............(*)"

Now , from the rank-nullity theorem ,

"rank(T)+nullity(T)=rank(ST)+nullity (ST)=n"

Now , from (*) we conclude that ,

"rank(ST)\leq rank(T)" .